Circular No. 9122

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 D3 (WISE) A. Mainzer, Jet Propulsion Laboratory, reports that an object dis-

covered on WISE satellite images (discovery observation tabulated below) appears clearly extended in 12-μm images with a ‘hazy’ coma of diameter ∼ 20′′ and a nondistinct core; in 22-μm images, the object appears elongated with size 20′′ × 30′′. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, G. J. Garradd reports that 0.5-m Uppsala Schmidt CCD images reveal a magnitude of 20.5 on Mar. 8.44 UT (comet slightly fuzzier than surrounding stars, but involved with a star).

2010 UT α_{2000} δ_{2000} Observer
Feb. 26.52274 6^{h}41^{m}09.^{s}98 −73^{o}24′46′′.8 WISE

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2010-E34.

T = 2010 Sept. 2.349 TT  ω = 304.^{o}535
Ω = 255.193 }2000.0
q = 4.23610 AU  i = 76.354

GK PERSEI E. O. Waagen and A. Henden, AAVSO, report that GK Per has risen

at least a half magnitude above its quiescent visual magnitude of ∼ 13.1. Waagen notes that, after many years of quite-similar-looking outbursts ∼ 1100–1300 days apart, rising to m_{v} ∼ 10.1 in 24–30 days, and lasting 53– 73 days, the last two outbursts were very different, with the Sept. 2004 outburst coming only 940 days after the preceding outburst, taking > 60 days to rise to a maximum of m_{v} = 10.6 and lasting ∼ 100 days, and the Dec. 2006 outburst coming about 810 days later, rising only to m_{v} = 11.5 with oscillations to mag 12.2 before returning to minimum ∼ 100 days after outburst. The current probable outburst is occurring ∼ 1180 days after the last outburst. Close monitoring is very important and is strongly urged. Selected recent visual magnitude estimates of GK Per reported to the AAVSO: Mar. 1.860 UT, 13.0 (G. Poyner, Birmingham, England); 4.805, 13.0 (P. Schmeer, Bischmisheim, Germany); 5.014, 13.0 (M. Komorous, London, Ontario, Canada); 6.090, 12.8 (J. Bortle, Stormville, NY, U.S.A.); 6.801, 12.7 (Schmeer); 7.030, 12.6 (Bortle); 7.795, 12.4 (Schmeer); 7.839, 12.3 (Poyner); 8.020, 12.5 (Komorous); 8.026, 12.4 (Bortle). Photoelectric V - band magnitudes from E. van Ballegoij, Heesch, The Netherlands: Mar. 4.917, 13.03; 6.870, 12.34.

2010 March 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green